OpenStudy (anonymous):

PLEASE HELP!!!!!!! An equation is shown below. 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (anonymous):

PART A) lets say x=c then y=9c+2 So all the possible solutions are (c,9c+2) (you can also keep x)

OpenStudy (anonymous):

Part B) For c=1,2,3 you have 3 different sets of solutions

OpenStudy (anonymous):

Part C) one equation would be y=9x

OpenStudy (anonymous):

a) solve for y, y=9x+2, now plug zero for x solve for y, y=9*(0)+2=2, so one solution is (0,2) and the other one....plug zero for y and solve for x 9x+2=0, x=-2/9 second solution is (-2/9,0) third solution is trivial case....to find inconsistent solution just find an equation that is parallel to this one..meaning they have the same slope so y=9x+5 would be inconsistent solution

OpenStudy (anonymous):

You guys r a lifesaver can you help me with one more question please

OpenStudy (anonymous):

A system of equations is shown below. 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Nevermind I got it